OpenStudy (anonymous):

find the critical numbers for :

OpenStudy (anonymous):

abs (x^2-1)

OpenStudy (amistre64):

what would you define as a critical number?

OpenStudy (amistre64):

if its about derivatives: the derivative of abs(u) = u' u/abs(u)

OpenStudy (amistre64):

you want zeros and undefineds

OpenStudy (anonymous):

that makes f prime (x) = 0

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

0, and -+1

OpenStudy (anonymous):

I know that , but how ?

OpenStudy (amistre64):

how what?

OpenStudy (anonymous):

how did you find these critical numbers

OpenStudy (amistre64):

\[|x^2-1|'=2x\frac{x^2-1}{|x^2-1|}\] set it equal to zero, and since its a fraction, a 0 denominator is undefined

OpenStudy (amistre64):

x=0, -1, 1

OpenStudy (amistre64):

2 conditions for critical values: f' = 0 f' is undefined people often forget about the second part

OpenStudy (anonymous):

ok so you took the numerator and set it equal to zero , then got x=1 and x=-1 ? since we can simplify it to (x-1)(x+1) ?

OpenStudy (anonymous):

and we got a zero from 2x

OpenStudy (amistre64):

in this case, thats a good way to see it. but actaully, i got a zero from 2x, and i got the undefined parts from the factors of x^2-1

OpenStudy (amistre64):

0/0 is undefined

OpenStudy (anonymous):

i know , but does it matter if I took the numerator in this case ? , if the numerator was x then yes I would look at a way to get it undefined instead of a 0.

OpenStudy (anonymous):

thanks by the way

OpenStudy (amistre64):

it does not matter in this case since the top and bottom are the same relative function. the absolute value parts acts like a switch that changes from 1 to -1 and back to 1 again. the only place that it is not defined at is when the top and bottom are 0

OpenStudy (anonymous):

thank you so much.

OpenStudy (amistre64):

youre welcome